### v2

@@ 3.3 Extensions
For clarity of exposition, suppose the regressors on the RHS of (2.1) include constant 1 and a scalar-valued $X_{it}$. It is straightforward to generalize the idea to cover $d_{X}$-dimensional $X_{it}$ for $d_{X}\geq 2$. With a slight abuse of notation, let $\left(\beta_{it},\omega_{it}\right)\in\mathbb{R}^{2}$ where $\beta_{it}$ is the random coefficient for $X_{it}$ and $\omega_{it}$ is for constant 1. Since the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$ is driven only by exogenous $Z_{it}$, (3.4) holds for any measurable function of $\beta_{it}$. Thus, I have
$\mathbb{E}\left[\left.\beta_{it}^{2}\right|X_{it},V_{it},W_{i}\right]$ $=\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right],$ (3.10)
and similarly for $\mathbb{E}\left[\left.\omega_{it}^{2}\right|X_{it},V_{it},W_{i}\right]$ and $\mathbb{E}\left[\left.\omega_{it}\beta_{it}\right|X_{it},V_{it},W_{i}\right]$. Taking the conditional expectation of both sides of (2.1) squared given $\left(X_{it},V_{it},W_{i}\right)$ yields
$\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]=X_{it}^{2}\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]+2X_{it}\mathbb{E}\left[\left.\beta_{it}\omega_{it}\right|V_{it},W_{i}\right]+\mathbb{E}\left[\left.\omega_{it}^{2}\right|V_{it},W_{i}\right].$ (3.11)
Then, I have
$\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]$ $=\left(\partial^{2}\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it}^{2}\right)/2,$ (3.12)
$\mathbb{E}\left[\left.\beta_{it}\omega_{it}\right|V_{it},W_{i}\right]$ $=\left(\partial\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it}-2X_{it}\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]\right)/2,\text{ and}$
$\mathbb{E}\left[\left.\omega_{it}^{2}\right|V_{it},W_{i}\right]$ $=\mathbb{E}\left[\left.Y_{it}^{2}\right|X_{it},V_{it},W_{i}\right]-X_{it}^{2}\mathbb{E}\left[\left.\beta_{it}^{2}\right|V_{it},W_{i}\right]-2X_{it}\mathbb{E}\left[\left.\beta_{it}\omega_{it}\right|V_{it},W_{i}\right],$
which identify the conditional (given $X_{it}$) and unconditional second-order moments of $\beta_{it}$ and $\omega_{it}$ by the LIE.
The above technique can also be used to identify the intertemporal correlations of the random coefficients. For example, $\mathbb{E}\left[\left.\beta_{it}\beta_{is}\right|X_{it},X_{is},V_{it},V_{is},W_{i}\right]$ can be identified from $\mathbb{E}\left[\left.Y_{it}Y_{is}\right|X_{it},X_{is},V_{it},V_{is},W_{i}\right]$ for any $t\neq s$ following an argument similar to (3.11)–(3.12).
@@
The identification argument of this paper goes through when I further include ex-post shocks $\upsilon_{it}$ to $\beta_{it}$ (i.e., $\beta_{it}=\beta\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$) and $\epsilon_{it}$ to $Y_{it}$ (i.e., $Y_{it}=X_{it}^{{}^{\prime}}\beta_{it}+\epsilon_{it}$), where both $\upsilon_{it}$ and $\epsilon_{it}$ are independent of all other variables with $\mathbb{E}\epsilon_{it}=0$. Note that due to the inclusion of $\upsilon_{it}$, $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ becomes a possibly time-varying function of $X_{it}$ as the distribution of $\upsilon_{it}$ may depend on time $t$. It does not affect my identification analysis because $\upsilon_{it}$ is independent of all other variables. I present the changes required to the proof due to the inclusion of $\upsilon_{it}$ into $\beta_{it}$ at the end of the proof of Theorem 1. I follow the more general specification $\beta_{it}=\beta\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$ in the estimation and examine its impact via simulations.
As for $\epsilon_{it}$, since it is independent of all other variables, has mean zero, and enters (2.1) in an additive way, equations (3.4)–(3.8) hold as before. Thus, the identification of $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ is not affected. However, it has an impact on the identification of higher-order moments of $\beta_{it}$. For example, in (3.12) $\mathbb{E}\left[\left.\omega_{it}^{2}\right|V_{it},W_{i}\right]$ is not identified if $\epsilon_{it}$ is included with $\mathbb{E}\epsilon_{it}^{2}>0$. One way to handle this issue is by following Arellano and Bonhomme (2012) to impose and exploit additional assumptions on how $\epsilon_{it}$ evolves over time (e.g., $\epsilon_{it}$ follows an ARMA(p,q) process).
Finally, when exogenous covariates $Z_{1,it}$ are available and included in $X_{it}$ [^20] If $Z_{1,it}$ is included in $X_{it}$, I can write $X_{it}=\left(U_{it}^{{}^{\prime}},Z_{1,it}^{{}^{\prime}}\right)^{{}^{\prime}}$, $U_{it}=g\left(Z_{it},A_{i},\eta_{it}\right)$, and $Z_{it}=\left(Z_{1,it}^{{}^{\prime}},Z_{2,it}^{{}^{\prime}}\right)^{{}^{\prime}}$., my identification analysis actually benefits from it because Assumptions 13 concern only endogenous covariates $U_{it}$ of $X_{it}$. Thus, I replace $X_{it}$ by $U_{it}$ in Assumptions 13 and in the construction of $V_{it}$. Furthermore, as discussed in footnote 16, the unconditional variation of $Z_{1,it}$ makes Assumption 4 and 4 easier to be satisfied. The details on how to adapt the current analysis to include $Z_{1,it}$ are provided at the end of the proof of Theorem 1.
@@ 4 Estimation and Large Sample Theory
In this section, I show how to estimate $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ via three-step series estimators based on the constructive identification argument. Next, I establish the convergence rates of the proposed estimators. Finally, I prove that the estimators are asymptotically normal and provide consistent estimators for their asymptotic covariance matrices.
@@ 4.1 Estimation
As discussed in Extension 3 of Section 3, I include the ex-post shock $\upsilon_{it}$ in $\beta_{it}$, resulting in $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]\eqqcolon b_{1t}\left(V_{it},W_{i}\right)$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]\eqqcolon b_{t}\left(X_{it}\right)$. The parameters of interest are
$\overline{b}\coloneqq\mathbb{E}\beta_{it}\ \text{and}\ b_{t}\left(x\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}=x\right].$ (4.1)
$\overline{b}$ is the APE over the unconditional distribution of $\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$. $b_{t}\left(x\right)$ is the LAR function for a subpopulation characterized by $X_{it}=x$ in period $t$. Both objects are useful for answering policy-related questions. For example, plugging realizations $x_{it}$ of $X_{it}$ into $b_{t}\left(x\right)$ provides a fine approximation to $\beta_{it}$ for agent $i$ in time $t$. Therefore, distributional properties of $\beta_{it}$ may be inferred from the distribution of $b_{t}\left(x_{it}\right)$.
I propose to estimate the parameters in (4.1) with three-step series estimators. For clarity of exposition, I let $d_{X}=1$, and highlight the changes required for $d_{X}>1$ when necessary. To fix ideas, I follow Liu, Poirier, and Shiu (2024) to let $W_{i}=T^{-1}\sum_{t=1}^{T}X_{it}$, which can be motivated by generalizing the method of Mundlak (1978).
First, for each $t$, I estimate $V_{t}\left(x,z,w\right)\coloneqq F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.x\right|z,w\right)$ by regressing $\mathbf{\mathbbm{1}}\left\{X_{it}\leq x\right\}$ on the basis functions $q^{M_{1}}\left(\cdot\right)$ of $\left(Z_{it},W_{i}\right)$ with trimming function $\tau\left(\cdot\right)$:
@@
$=\tau\left(q^{M_{1}}\left(z,w\right)^{{}^{\prime}}n^{-1}\widehat{Q}_{t}^{-1}\sum_{j=1}^{n}q_{jt}\mathbf{\mathbbm{1}}\left\{x_{jt}\leq x\right\}\right)$
$\eqqcolon\tau\left(q^{M_{1}}\left(z,w\right)^{{}^{\prime}}\widehat{\gamma}_{t}^{M_{1}}\left(x\right)\right),$ (4.2)
[…]
Next, define $S\coloneqq\left(X,V,W\right)$ and let $\mathcal{X},\ \mathcal{Z},\ \mathcal{V},\ \mathcal{W},$ and $\mathcal{S}$ denote the supports of $X,\ Z,\ V,\ W,$ and $S$, respectively. Let $V_{it}\coloneqq V_{t}\left(X_{it},Z_{it},W_{i}\right),$$\widehat{V}_{it}\coloneqq\widehat{V}_{t}\left(X_{it},Z_{it},W_{i}\right)$, and $\widehat{v}_{it}\coloneqq\widehat{V}_{t}\left(x_{it},z_{it},w_{i}\right)$. For any $s=\left(x,v,w\right)\in\mathcal{S}$, I estimate $G_{t}\left(s\right)\coloneqq\mathbb{E}\left[\left.Y_{it}\right|S_{it}=s\right]$ by regressing $Y_{it}$ on the basis functions $p^{M_{2}}\left(\cdot\right)$ of $\left(X_{it},\widehat{V}_{it},W_{i}\right)$:
$\widehat{G}_{t}\left(s\right)=p^{M_{2}}\left(s\right)^{{}^{\prime}}n^{-1}\widehat{P}_{t}^{-1}\widehat{p}_{t}^{{}^{\prime}}y_{t}\eqqcolon p^{M_{2}}\left(s\right)^{{}^{\prime}}\widehat{\alpha}_{t}^{M_{2}},$
where $\widehat{P}_{t}:=n^{-1}\sum_{i=1}^{n}\widehat{p}_{it}\widehat{p}_{it}^{{}^{\prime}}$, $\widehat{p}_{it}:=p^{M_{2}}\left(x_{it},\widehat{v}_{it},w_{i}\right)$, $\widehat{p}_{t}:=\left(\widehat{p}_{1t},...,\widehat{p}_{nt}\right)^{{}^{\prime}}$, and $y_{t}:=\left(y_{1t},...,y_{nt}\right)^{{}^{\prime}}$. Following Newey, Powell, and Vella (1999), I let $p^{M_{2}}\left(s\right)=x\otimes p^{m_{2}}\left(v,w\right)$ by exploiting the index structure of (2.1), which enables a faster convergence rate of $\widehat{G}_{t}\left(s\right)$ to $G_{t}\left(s\right)$.
@@
$\widehat{b}_{1t}\left(v,w\right)=\left(\widehat{\mathbb{E}}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|\widehat{V}_{it}=v,W_{i}=w\right]\right)^{-1}\widehat{\mathbb{E}}\left[\left.X_{it}Y_{it}\right|\widehat{V}_{it}=v,W_{i}=w\right],$ (4.3)
@@
where $\widehat{\mathbb{E}}\left[\left.X_{it}Y_{it}\right|\widehat{V}_{it},W_{i}\right]$ is obtained by regressing each coordinate of $X_{it}Y_{it}$ on the basis functions of $\left(\widehat{V}_{it},W_{i}\right)$ and similarly for $\widehat{\mathbb{E}}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|\widehat{V}_{it},W_{i}\right]$. When (3.7) is used, I estimate $b_{1t}\left(v,w\right)$ by
@@
$\widehat{b}_{1t}\left(v,w\right)=\partial\widehat{G}_{t}\left(s\right)/\partial x=\left(I_{d_{X}}\otimes p^{m_{2}}\left(v,w\right)\right)^{{}^{\prime}}\widehat{\alpha}_{t}^{M_{2}}\eqqcolon\overline{p}^{M_{2}}\left(s\right)^{{}^{\prime}}\widehat{\alpha}_{t}^{M_{2}},$ (4.4)
@@
where the second equality holds by the chain rule. I follow (4.4) in what follows for its simplicity in deriving the asymptotic properties.
@@
To estimate $\overline{b}$ and $b_{t}\left(x\right)$, by the LIE I regress $\widehat{b}_{1t}\left(\widehat{V}_{it},W_{i}\right)$ on the basis function $r^{M_{3}}\left(\cdot\right)$ of constant one and $X_{it}$, respectively:
@@
$\widehat{\overline{b}}$ $=\left(nT\right)^{-1}\sum_{t=1}^{T}\sum_{i=1}^{n}\widehat{b}_{1t}\left(\widehat{v}_{it},w_{i}\right),\text{ and}$
$\widehat{b}_{t}\left(x\right)$ $=r^{M_{3}}\left(x\right)^{{}^{\prime}}n^{-1}\widehat{R}_{t}^{-1}r_{t}^{{}^{\prime}}\widehat{B}_{t}\eqqcolon r^{M_{3}}\left(x\right)^{{}^{\prime}}\widehat{\rho}_{t}^{M_{3}},$ (4.5)
where $\widehat{R}_{t}:=n^{-1}\sum_{i=1}^{n}r_{it}r_{it}^{{}^{\prime}}$, $r_{it}:=r^{M_{3}}\left(x_{it}\right)$, $r_{t}:=\left(r_{1t},...,r_{nt}\right)^{{}^{\prime}}$, and $\widehat{B}_{t}:=\left(\widehat{b}_{1t}\left(\widehat{v}_{1t},w_{1}\right),...,\widehat{b}_{1t}\left(\widehat{v}_{nt},w_{n}\right)\right)^{{}^{\prime}}$.

### v3

@@ 3.3 Extensions
I identify the $P^{\text{th}}$-order moments of $\beta_{it}$. For clarity of illustration, suppose the regressor is $X_{it}=(K_{it},L_{it})^{\prime}$ and $\beta_{it}=(\beta_{it,K},\beta_{it,L})^{\prime}$. The extension to $d_{X}>2$ is straightforward. Since the residual variation in $X_{it}$ given $(V_{it},W_{i})$ is driven only by exogenous $Z_{it}$, (3.9) holds for any measurable function $h$ of $\beta_{it}$:
$\mathbb{E}\left[h(\beta_{it})\mid X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[h(\beta_{it})\mid V_{it},W_{i}\right].$ (3.20)
In particular, let $h=(h_{1},\ldots,h_{P})^{\prime}$ and $h_{q}(\beta_{it})=\beta_{it,K}^{q}\beta_{it,L}^{P-q}$ for $q=0,\ldots,P$. By (1.1), the conditional $P^{\text{th}}$-order moment of $Y_{it}$ given $(X_{it},V_{it},W_{i})$ is
$\mathbb{E}\left[Y_{it}^{P}\mid X_{it},V_{it},W_{i}\right]=\sum_{q=0}^{P}\binom{P}{q}K_{it}^{q}L_{it}^{P-q}\mathbb{E}\left[\beta_{it,K}^{q}\beta_{it,L}^{P-q}\mid V_{it},W_{i}\right].$ (3.21)
Under $P$-times differentiability of $\mathbb{E}\left[Y_{it}^{P}\mid X_{it},V_{it},W_{i}\right]$ in $X_{it}$, the coefficients on $K_{it}^{P}$ and $L_{it}^{P}$ are identified via differentiation, yielding
$\mathbb{E}\left[\beta_{it,K}^{q}\beta_{it,L}^{P-q}\mid V_{it},W_{i}\right]$ $=\dfrac{1}{P!}\frac{\partial^{P}}{\partial k^{q}\partial l^{P-q}}\mathbb{E}\left[Y_{it}^{P}\mid K_{it}=k,L_{it}=l,V_{it},W_{i}\right],\text{ for }q=0,\ldots,P,$ (3.22)
and hence the $P^{\text{th}}$-order moments of $\beta_{it}$ follow by the law of iterated expectations
$\mathbb{E}[\beta_{it,K}^{q}\beta_{it,L}^{P-q}]$ $=\mathbb{E}\left[\dfrac{1}{P!}\frac{\partial^{P}}{\partial k^{q}\partial l^{P-q}}\mathbb{E}\left[Y_{it}^{P}\mid K_{it}=k,L_{it}=l,V_{it},W_{i}\right]\right],\text{ for }q=0,\ldots,P.$ (3.23)
Repeating this argument for all $P\in\mathbb{N}$ identifies the entire moment sequence of $\beta_{it}$. Under standard moment-determinacy conditions (e.g., Stoyanov (2000)), this sequence uniquely determines the distribution of $\beta_{it}$.
@@
The identification argument goes through when I include ex-post shocks $\upsilon_{it}$ to $\beta_{it}$ (i.e., $\beta_{it}=\beta\left(A_{i},\varepsilon_{it},\upsilon_{it}\right)$) and $\widetilde{\epsilon}_{it}$ to $Y_{it}$ (i.e., $Y_{it}=X_{it}^{\prime}\beta_{it}+\widetilde{\epsilon}_{it}$ with $\mathbb{E}\widetilde{\epsilon}_{it}=0$), where $\upsilon_{it}$ and $\widetilde{\epsilon}_{it}$ are independent of all other variables. Note that $\beta(\cdot)$ may a priori be time-varying and the joint distribution of $(\varepsilon_{it},\eta_{it})$ may also vary with $t$. Introducing ex-post shocks $\upsilon_{it}$ provides an additional and economically distinct source of time variation (e.g., unexpected technology shock). While the identification result remains valid with the inclusion of $\upsilon_{it}$, I present the required changes to the proof at the end of the proof of Theorem 1. I also include $\upsilon_{it}$ into $\beta_{it}$ when estimating the APE and LAR and examine its impact via simulations. As for $\widetilde{\epsilon}_{it}$, since it is independent of all other variables and enters (2.1) in an additive way, equations (3.9)–(3.13) hold as before. Thus, the identification analysis is unaffected.
The presence of exogenous covariates in $X_{it}$ can strengthen identification. The key point is that Assumptions 13 pertain only to the endogenous covariates of $X_{it}$. In particular, the control variables $V_{it}$ and $W_{i}$ depend only on the endogenous covariates of $X_{it}$, which alleviates concerns that the controls may be high-dimensional and makes the required residual variation in $X_{it}$ easier to satisfy. Moreover, as discussed in Appendix A.3, unconditional variation in exogenous covariates of $X_{it}$ can be directly exploited to make Assumptions 4 and 4 easier to satisfy. Details on how to adapt the analysis to incorporate exogenous covariates are provided at the end of the proof of Theorem 1.
@@ 4 Estimation and Large Sample Theory
In this section, I first describe how to estimate the parameters using three-step series estimators. I then establish the convergence rates of the proposed estimators, and finally show that they are asymptotically normal.
@@ 4.1 Estimation
The parameters of interest are
$\overline{b}=T^{-1}\sum_{t=1}^{T}\mathbb{E}\beta_{it}\quad\text{and}\quad b_{t}\left(x\right)\coloneqq\mathbb{E}\left[\left.\beta_{it}\right|X_{it}=x\right].$ (4.1)
$\overline{b}$ is the pooled APE across all firms and periods. $b_{t}\left(x\right)$ is the LAR function for a subpopulation characterized by $X_{it}=x$ in period $t$ and is also useful for answering policy-related questions. For example, plugging realizations $x_{it}$ of $X_{it}$ into $b_{t}\left(x\right)$ provides a fine approximation to $\beta_{it}$.
I propose to estimate the parameters in (4.1) with three-step series estimators. For clarity of exposition, I set $d_{X}=1$ and highlight the modifications required for $d_{X}>1$ when needed. To fix ideas, I follow Liu, Poirier, and Shiu (2025) to let $W_{i}=T^{-1}\sum_{t=1}^{T}X_{it}$, which can be motivated by generalizing the method of Mundlak (1978).
First, for each $t$, I estimate $V_{t}\left(x,z,w\right)\coloneqq F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.x\right|z,w\right)$ by regressing $\mathbf{\mathbbm{1}}\left\{X_{it}\leq x\right\}$ on the basis functions $q^{M_{1}}$ of $\left(Z_{it},W_{i}\right)$ with trimming function $\tau$:
@@
$=\tau\left(q^{M_{1}}\left(z,w\right)^{\prime}n^{-1}\widehat{Q}_{t}^{-1}\sum_{j=1}^{n}q_{jt}\mathbf{\mathbbm{1}}\left\{x_{jt}\leq x\right\}\right)$
$\eqqcolon\tau\left(q^{M_{1}}\left(z,w\right)^{\prime}\widehat{\gamma}_{t}^{M_{1}}\left(x\right)\right),$ (4.2)
where $\widehat{Q}_{t}:=n^{-1}\sum_{i=1}^{n}q_{it}q_{it}^{\prime}$ and $q_{it}:=q^{M_{1}}\left(z_{it},w_{i}\right)$. Examples of $q^{M_{1}}$ include power series and spline functions. When $d_{X}>1$, I regress $\mathbf{\mathbbm{1}}\left\{X_{it,l}\leq x_{l}\right\}$ on the basis functions $q^{M_{1}}$ of $\left(Z_{it},W_{i}\right)$ with trimming function $\tau$ for each $l$ and obtain $\widehat{V}_{t}\left(x,z,w\right)=\left(\widehat{V}_{t,1}\left(x,z,w\right),\ldots,\widehat{V}_{t,d_{X}}\left(x,z,w\right)\right)^{\prime}$. I highlight two properties of $\widehat{V}_{t}\left(x,z,w\right)$. First, the regression coefficient $\widehat{\gamma}_{t}^{M_{1}}\left(x\right)$ in (4.2) depends on $x$ because the dependent variable $\mathbf{\mathbbm{1}}\left\{X_{it}\leq x\right\}$ is a function of $x$. This fact causes the convergence rate of $\widehat{V}_{t}\left(x,z,w\right)$ to be slower than the standard rates for series estimators (Imbens and Newey (2009)). Second, a trimming function $\tau$ is needed since I am estimating a conditional CDF. An example of $\tau$ is $\tau\left(x\right)=\mathbf{\mathbbm{1}}\left\{x>0\right\}\cdot\min\left(x,1\right)$.
Next, define $S\coloneqq\left(X,V,W\right)$ and let $\mathcal{X},\ \mathcal{Z},\ \mathcal{V},\ \mathcal{W},$ and $\mathcal{S}$ denote the supports of X,Z,V,W,X,\ Z,\ V,\ W, and $S$, respectively. Let $V_{it}\coloneqq V_{t}\left(X_{it},Z_{it},W_{i}\right),$$\widehat{V}_{it}\coloneqq\widehat{V}_{t}\left(X_{it},Z_{it},W_{i}\right)$, and $\widehat{v}_{it}\coloneqq\widehat{V}_{t}\left(x_{it},z_{it},w_{i}\right)$. For any $s=\left(x,v,w\right)\in\mathcal{S}$, I estimate $G_{t}\left(s\right)\coloneqq\mathbb{E}\left[\left.Y_{it}\right|S_{it}=s\right]$ by regressing $Y_{it}$ on the basis functions $p^{M_{2}}$ of $\left(X_{it},\widehat{V}_{it},W_{i}\right)$:
$\widehat{G}_{t}\left(s\right)=p^{M_{2}}\left(s\right)^{\prime}n^{-1}\widehat{P}_{t}^{-1}\widehat{p}_{t}^{\prime}y_{t}\eqqcolon p^{M_{2}}\left(s\right)^{\prime}\widehat{\alpha}_{t}^{M_{2}},$ (4.3)
where $\widehat{P}_{t}:=n^{-1}\sum_{i=1}^{n}\widehat{p}_{it}\widehat{p}_{it}^{\prime}$, $\widehat{p}_{it}:=p^{M_{2}}\left(x_{it},\widehat{v}_{it},w_{i}\right)$, $\widehat{p}_{t}$ is an $n\times M_{2}$ matrix with $\widehat{p}_{it}^{\prime}$ being its $i^{\text{th}}$ row, and $y_{t}$ is the vector of $y_{it}$s. Following Newey, Powell, and Vella (1999), I let
$p^{M_{2}}\left(s\right)=x\otimes p^{m_{2}}\left(v,w\right)$ (4.4)
by exploiting the index structure of (2.1). Hence, the effective degree that matters for the convergence rate is $m_{2}$ since $M_{2}=d_{X}\times m_{2}$ and $d_{X}$ is finite.
@@
$\widetilde{b}_{1t}\left(v,w\right)=\left(\widehat{\mathbb{E}}\left[\left.X_{it}X_{it}^{\prime}\right|\widehat{V}_{it}=v,W_{i}=w\right]\right)^{-1}\widehat{\mathbb{E}}\left[\left.X_{it}Y_{it}\right|\widehat{V}_{it}=v,W_{i}=w\right],$ (4.5)
@@
where $\widehat{\mathbb{E}}\left[\left.X_{it}Y_{it}\right|\widehat{V}_{it},W_{i}\right]$ is obtained by regressing each coordinate of $X_{it}Y_{it}$ on the basis functions of $\left(\widehat{V}_{it},W_{i}\right)$ and similarly for $\widehat{\mathbb{E}}\left[\left.X_{it}X_{it}^{\prime}\right|\widehat{V}_{it},W_{i}\right]$. When (3.12) is used, I estimate $b_{1t}\left(v,w\right)$ by
@@
$\widehat{b}_{1t}\left(v,w\right)=\partial\widehat{G}_{t}\left(s\right)/\partial x=\left(I_{d_{X}}\otimes p^{m_{2}}\left(v,w\right)\right)^{\prime}\widehat{\alpha}_{t}^{M_{2}}\eqqcolon\overline{p}^{M_{2}}\left(s\right)^{\prime}\widehat{\alpha}_{t}^{M_{2}},$ (4.6)
@@
where the second equality holds by the chain rule. I follow (4.6) in what follows, as it simplifies both implementation and the derivation of asymptotic properties.
Remark 2 (Compare $\boldsymbol{\widetilde{b}_{1t}\left(v,w\right)}$ with $\boldsymbol{\widehat{b}_{1t}\left(v,w\right)}$).
While both approaches are valid, this paper recommends using the derivative-based estimator $\widehat{b}_{1t}(v,w)$ rather than the inversion-based estimator $\widetilde{b}_{1t}(v,w)$ in practice when Assumption 4 holds. The derivative-based route operates through the estimated one-dimensional regression function $\widehat{G}_{t}(x,v,w)$ and obtains $\widehat{b}_{1t}(v,w)$ via the analytic partial derivative with respect to $x$. This approach exploits the linear random-coefficient structure directly, avoids an additional “matrix estimation + inversion” step, and leads to a cleaner and more stable asymptotic argument in the paper’s framework. By contrast, $\widetilde{b}_{1t}(v,w)$ requires estimating and inverting the conditional second-moment matrix $\mathbb{E}[X_{it}X_{it}^{\prime}\mid V_{it}=v,W_{i}=w]$, whose sample analogue can be ill-conditioned in finite samples and hence numerically unstable (Carrasco, Florens, and Renault (2007)). This inversion also complicates the asymptotic analysis, because one must control how estimation error in the conditional second moment propagates through a matrix inverse.
@@
To estimate $\overline{b}$ and $b_{t}\left(x\right)$, by the law of iterated expectations I regress $\widehat{b}_{1t}\left(\widehat{V}_{it},W_{i}\right)$ on the basis function $r^{M_{3}}$ of constant one and $X_{it}$, respectively:
@@
$\widehat{\overline{b}}=\dfrac{1}{nT}\sum_{t=1}^{T}\sum_{i=1}^{n}\widehat{b}_{1t}\left(\widehat{v}_{it},w_{i}\right)\quad\text{and}\quad\widehat{b}_{t}\left(x\right)=r^{M_{3}}\left(x\right)^{\prime}n^{-1}\widehat{R}_{t}^{-1}r_{t}^{\prime}\widehat{B}_{t}\eqqcolon r^{M_{3}}\left(x\right)^{\prime}\widehat{\rho}_{t}^{M_{3}},$ (4.7)
where $\widehat{R}_{t}:=n^{-1}\sum_{i=1}^{n}r_{it}r_{it}^{\prime}$, $r_{it}:=r^{M_{3}}(x_{it})$, $r_{t}$ is an $n\times M_{3}$ matrix with $r_{it}^{\prime}$ being its $i^{\text{th}}$ row, and $\widehat{B}_{t}$ is an $n\times d_{X}$ matrix with $\widehat{b}_{1t}(\widehat{v}_{it},w_{i})^{\prime}$ being its $i^{\text{th}}$ row.
Remark 3 (Computational Considerations).
While the three-step estimators are based on standard series regressions and are typically computationally manageable, two implementation aspects are worth noting for large datasets. First, when estimating $V_{t}\left(x,z,w\right)$, a key computational simplification is that the regressors—the basis functions $q^{M_{1}}(Z_{it},W_{i})$—are identical across all thresholds, with only the dependent variable $\mathbf{\mathbbm{1}}\{X_{it}\leq x\}$ changing in $x$. Hence, I can compute the inverse Gram matrix $\widehat{Q}_{t}^{-1}$ once and reuse it to obtain the coefficient vector $\widehat{\gamma}_{t}^{M_{1}}\left(x\right)$ for every threshold $x_{it}$. This reuse is substantially cheaper than running $n$ fully separate regressions. Second, when estimating $G_{t}\left(s\right)$, I exploit the linear structure in $X_{it}$ using the separable specification $p^{M_{2}}\left(s\right)=x\otimes p^{m_{2}}\left(v,w\right)$, which reduces the effective regressor dimensionality and alleviates the associated computational burden.